# Approach to stationarity for the KPZ fixed point with boundaries

### v1

@@ 1 Brownian excursions
2 Non-intersecting Wiener processes and Brownian excursions
@@
3 Expectation values of Brownian averages
@@
[…]
@@ V Bethe ansatz
1 TASEP with periodic boundaries
@@
2 Open TASEP with $\alpha=\beta=1$
a Analytic continuations in the variable $B$
$y_{0}^{+}$$y_{0}^{-}$$y_{1}^{+}$$y_{1}^{-}$$y_{2}^{+}$$y_{2}^{-}$$y_{3}^{+}$$y_{3}^{-}$$y_{4}^{+}$$y_{4}^{-}$
Figure S9: Image of the slit plane $\mathbb{C}\setminus[0,\infty)$ by the Bethe root functions $y_{j}^{\pm}(B)$ for open TASEP with $L=3$ and $\alpha=\beta=1$.
@@ 2 Open TASEP with $\alpha=\beta=1$
We note that the image of the branch cut $w_{j}\in(-\infty,0]$ by $y_{j}^{\pm}=w_{j}\pm w_{j}\sqrt{1-w_{j}^{-2}}$ is the unit circle $|y_{j}^{\pm}|=1$, and $|y_{j}^{+}|>1$, $|y_{j}^{-}|<1$ outside the cut. Labelling the $y_{j}^{\pm}$ as in figure S9, we observe that analytic continuation across the cuts $(B_{*},\infty)$ and $(0,B_{*})$ is realized by the action on the set $\{j_{\theta_{j}},j=0,\ldots,L+1\}$ of the respective permutations (written as a product of two cycles) $\mathfrak{a}=(0_{+},0_{-})\,(1_{+},2_{+},\ldots,(L+1)_{+},1_{-},2_{-},\ldots,(L+1)_{-})$ and $\mathfrak{b}=(0_{+},1_{+},\ldots,(L+1)_{+})\,(0_{-},1_{-},\ldots,(L+1)_{-})$.
The condition $\prod_{j=0}^{L+1}\theta_{j}=(-1)^{L}$ for acceptable choices of the Bethe roots $y_{j}=y_{j}^{\theta_{j}}$ is in particular preserved by $\mathfrak{a}$ and $\mathfrak{b}$. Similarly to TASEP with periodic boundaries, the Riemann surface $\mathcal{R}$ associated with the spectral curve $\det(\lambda\mathrm{Id}-M_{i}(\gamma))=0$ for open TASEP, independent of $i$ since $M_{i}(\gamma)$ and $M_{0}(\gamma)$ are related by a similarity transformation, is built from sheets ^$\widehat{\mathbb{C}}_{\Theta}$ for all $\Theta=(\theta_{0},\ldots,\theta_{L+1})$ verifying the condition $\prod_{j=0}^{L+1}\theta_{j}=(-1)^{L}$. The $2^{L+1}$ sheets are glued together in $\mathcal{R}$ according the action of the permutations $\mathfrak{a}$ and $\mathfrak{b}$ on $\Theta$.
b Differential $\omega$
@@
c Computation of $\kappa_{\rm st}$ at large $L$
@@
The stationary eigenstate corresponds to choosing the Bethe roots $y_{j}=y_{j}^{-}=w_{j}-w_{j}\sqrt{1-w_{j}^{-2}}$, which vanish when $B\to 0$ (corresponding to $w_{j}\to\infty$), unlike the $w_{j}^{+}$. Then, following [16, 29], one can write $\kappa$ for the stationary eigenstate at small enough $|B|$ as the contour integral around 00 with positive orientation
@@
with $\chi$ as in (16) and $v<1$, as stated in the letter.
d Computation of $\mathcal{N}_{\rm st}$ at large $L$

### v2

@@ IV.1 Brownian excursions
IV.2 Non-intersecting Wiener processes and Brownian excursions
@@
IV.3 Expectation values of Brownian averages
@@
$p_{y}(t,x)=\frac{\mathbb{P}(w>y,w(t)\in[x,x+\mathrm{d}x])}{\mathrm{d}x}=\frac{\mathrm{e}^{-\frac{x^{2}}{2t}}}{\sqrt{2\pi t}}-\frac{\mathrm{e}^{-\frac{(2y-x)^{2}}{2t}}}{\sqrt{2\pi t}}\propto y^{2}\;\text{when $y\to 0$}\;.$ (S83)
@@ V Bethe ansatz
V.1 TASEP with periodic boundaries
@@
V.2 Open TASEP with $\alpha=\beta=1$
V.2.a Analytic continuations in the variable $B$
@@ V.2 Open TASEP with $\alpha=\beta=1$
Similarly to TASEP with periodic boundaries, the Riemann surface $\mathcal{R}$ associated with the spectral curve $\det(\lambda\mathrm{Id}-M_{i}(\gamma))=0$ for open TASEP, independent of $i$ since $M_{i}(\gamma)$ and $M_{0}(\gamma)$ are related by a similarity transformation, is built from sheets ^$\widehat{\mathbb{C}}_{\Theta}$ for all $\Theta=(\theta_{0},\ldots,\theta_{L+1})$ verifying the condition $\prod_{j=0}^{L+1}\theta_{j}=(-1)^{L}$. The $2^{L+1}$ sheets are glued together in $\mathcal{R}$ according the action of the permutations $\mathfrak{a}$ and $\mathfrak{b}$ on $\Theta$.
V.2.b Differential $\omega$
@@
V.2.c Computation of $\kappa_{\rm st}$ at large $L$
@@
The stationary eigenstate corresponds to choosing the Bethe roots $y_{j}=y_{j}^{-}=w_{j}-w_{j}\sqrt{1-w_{j}^{-2}}$, which vanish when $B\to 0$ (corresponding to $w_{j}\to\infty$), unlike the $y_{j}^{+}$. Then, following [16, 31], one can write $\kappa$ for the stationary eigenstate at small enough $|B|$ as the contour integral around 00 with positive orientation
@@
with $\chi$ as in (S42) and $v<1$, as stated in the letter.
V.2.d Computation of $\mathcal{N}_{\rm st}$ at large $L$